Shift Summary Owl, August 5, 1999

Jim Hirschauer (jhirsch@cuhep.Colorado.EDU)
Thu, 5 Aug 1999 06:06:44 -0600 (MDT)

Shift Summary for owl shift, 5/Aug/99

On Shift: Arun Tripathi, Jim Hirschauer

Hours of beam delivered at all intensities: 4.0
Hours of beam delivered at nominal intensity: 4.0
Hours of E832 data (special runs not included): 3hrs 10 min
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: 50 min

In order to clear DYC busy's etc. (not cured by kfrend), initializing the
pipeline, stopping and starting runs etc.

========================================================================
Runs ended this shift: 14106, 14107
Run 14106 - 14 M - BEAM_832
Run 14107 - 6.9 M - BEAM_832
Run 14109 - in progress - BEAM_832

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Beam returns at 2:30. Smooth running until 5:00 when a whole mess of
problems results in the loss of 50 min of data-taking.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14106 1:30 Run ended due to lack of beam and desire to
incorporate new KE3 prescale.

14107 40 2:27 Beam returns

140 5:04 HCC inefficient alarm for several spills in a row.
Tried reloading ETOT, button came back red. All
hell breaks loose:
several DYC's busy, KFREND's do not fix. stop the
run and do drt_fresh_start for new run 14108.
after fresh start pipeline cr06 button came up
red (Error from command 39). rebooted the cr06
from the experts menu (with Julie's help),
KFRENDed and restarted the
run; found one quadrant missing in the HCC illum
plot. tried loading ET and KFRENDing, but no
triggers. several DYC and HCC busies were
enabled. called Julie and Rick. stopped run 14108
and did brute force init for everything. run
14109 seems ok ... for now

14109 1-50 6-7 intermittent aborts for some spills


========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

o Hot wire in kumquat 3Y-6 is being monitored and is not causing any
problems w/ trigger inefficiencies. There is no need to page the
expert unless frowny faces in "Kumquat Readout Errors" plot turn to
dead faces.